# TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning

### v1

@@ Abstract
Instruction tuning is essential for aligning large language models (LLMs) to downstream tasks and commonly relies on large, diverse corpora. However, small, high-quality subsets, known as coresets, can deliver comparable or superior results, though curating them remains challenging. Existing methods often rely on coarse, sample-level signals like gradients, an approach that is computationally expensive and overlooks fine-grained features. To address this, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention), a forward-only, token-centric framework. Instead of using gradients, TRIM operates by matching underlying representational patterns identified via attention-based "fingerprints" from a handful of target samples. Such an approach makes TRIM highly efficient and uniquely sensitive to the structural features that define a task. Coresets selected by our method consistently outperform state-of-the-art baselines by up to 9% on downstream tasks and even surpass the performance of full-data fine-tuning in some settings. By avoiding expensive backward passes, TRIM achieves this at a fraction of the computational cost. These findings establish TRIM as a scalable and efficient alternative for building high-quality instruction-tuning datasets.
TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning
Manish Nagaraj, Sakshi Choudhary, Utkarsh Saxena, Deepak Ravikumar, Kaushik Roy Department of Electrical and Computer Engineering, Purdue University {mnagara, choudh23, saxenau, dravikum, kaushik}@purdue.edu
@@ 1 Introduction
Large language models (LLMs) have become the de facto standard for a wide range of tasks, but their full potential is unlocked only after a critical, computationally expensive step: instruction tuning. While large, diverse instruction-tuning corpora have shown to be effective (Ouyang et al., 2022; Chung et al., 2024), a fundamental challenge persists: not all data is created equal. Recent findings have demonstrated that carefully curating a small, high-quality subset, or a “coreset”, can not only match but even surpass the performance of finetuning on the full dataset, all while using a fraction of the compute (Xia et al., 2024; Zhang et al., 2025a).
Current solutions often use influence-based methods to estimate sample-level data importance, relying on either gradients (Xia et al., 2024; Zhang et al., 2025b, c) or Hessians (Kwon et al., 2024; San Joaquin et al., 2024; Lin et al., 2024a). While effective, these are prohibitively expensive for large-scale data selection, requiring multiple backward passes, extensive memory, or a batch size of one to obtain per-sample gradients Pruthi et al. (2020). Representation similarity metrics (Hanawa et al., 2021; Zhang et al., 2018) alleviate the computational burden but still operate at the sample level. This coarse granularity introduces two key issues. First, a length bias, where deriving sample-wise scores from aggregated token-level metrics (e.g., loss, gradient magnitudes) creates a dependency on sequence length. Moreover, recent work shows that gradient norms in instruction-tuning samples are negatively correlated with sequence length. Xia et al. (2024). Second, a loss-centric bias, the next-token prediction objective distributes credit and blame uniformly across all tokens. In practice, however, softmax attention is inherently sparse, with performance often determined by a small subset of influential tokens Zhang et al. (2023); Ge et al. (2023); Lin et al. (2024b). Token-agnostic importance measures based solely on loss or gradients can misrank samples whose key signals are sparse and highly localized at a few tokens.
@@
In principle, the most faithful measure of influence would come from token-wise gradients (or Hessians), but this is computationally infeasible in practice. We therefore aim to develop efficient proxies that preserve token-level resolution while avoiding the prohibitive cost of gradient computation. This leads us to a central question: “Given only a few target samples for a downstream task, how do we efficiently identify high-impact instruction data for task-specific fine-tuning?"
@@
Figure 1: Overview of TRIM. (a) From multi-layer attention, we derive an aggregated token saliency signal that combines row (allocation sharpness) and column (received attention) signals. (b) Using a small target set, we compute one fingerprint per token class by saliency-weighted averaging of last-layer hidden states, capturing task-defining patterns (Section˜3.1). (c) For each candidate sample, token states are matched to their class fingerprints (cosine similarity), and token scores are pooled into a single relevance score for ranking (Section˜3.2); top-ranked samples form the instruction-tuning coreset.
@@
We propose a shift from the traditional sample-level, loss-centric paradigm to a token-level, representation-centric one, where we leverage the model’s hidden states, rich numerical vectors that capture a token’s meaning in context. Our core insight is that a sample’s true utility often resides not in the entire sequence but in a few informative tokens. To that effect, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention), a novel method that identifies influential data by matching the underlying attention patterns of a target task.
@@
As shown in Figure˜1, TRIM is a two-stage pipeline requiring only forward passes. Stage I constructs lightweight token fingerprints that capture the essence of a target task from a small set of samples. We begin by computing an aggregated token saliency score that combines two complementary signals: row saliency, which measures how sharply a token distributes its own attention, and column saliency, which measures how much attention it receives from other tokens (panel a). For each token class, we then compute a fingerprint as the saliency-weighted mean of the last-layer hidden states of all target-set tokens in that class (panel b), yielding a concise summary of what that class looks like to the model. Stage II uses these fingerprints to score samples in the large candidate corpus. For each sample, we compare each token’s last-layer hidden state to its class fingerprint via cosine similarity (panel c), then pool the resulting token-level scores into a single relevance score that reflects the sample’s alignment with the target task. Samples with the highest scores are selected to form the final coreset, comprising training examples whose token-level representations most strongly align with the target task’s structural patterns.
This design yields three advantages. First, TRIM delivers superior performance and efficiency: with only 5–10 samples from target task, it surpasses state-of-the-art methods (Xia et al., 2024; Zhang et al., 2025b; Yang et al., 2024) while running orders of magnitude faster, owing to its forward-only design that avoids per-sample gradient computation (Sections˜4.1 and 5). Second, it offers high structural fidelity by matching core representational patterns (e.g., syntax, mathematical operators), which benefits challenging downstream tasks (Section˜4.2). Finally, its token-level scoring inherently mitigates length bias, a common failure mode of sample-level methods (Section˜4.4).
Our contributions are:
  • A Token-Centric Framework. We introduce TRIM, to our knowledge the first token-level, representation-centric data-selection method that leverages contextualized token representations via attention-derived aggregated saliency.
@@
  • A Scalable and Efficient Algorithm. TRIM is a forward-only algorithm that is orders of magnitude faster than gradient-based alternatives, enabling practical and efficient data selection over massive instruction corpora.
@@
  • State-of-the-art empirical performance. With only a 5% coreset and 5-10 target samples, TRIM outperforms state-of-the-art methods by up to 9% on downstream reasoning tasks, and even exceeds full-data fine-tuning on some tasks.
@@ 2 Related Work
The challenge of curating coresets for instruction tuning has motivated a wide body of work on automated data selection (Cao et al., 2024; Chen et al., 2024; Zhou et al., 2023; Ding et al., 2023). Existing methods can be grouped into several paradigms. Influence-Based Selection methods estimate data importance by quantifying the causal effect of a candidate example on a target task. As exact influence functions (Koh and Liang, 2017) are infeasible for LLMs, recent works use scalable approximations. Some methods leverage Hessians Kwon et al. (2024); San Joaquin et al. (2024); Lin et al. (2024a), while other first-order alternatives track gradient similarity across checkpoints, as exemplified by the state-of-the-art LESS method (Xia et al., 2024). Forward-only approaches like CLD (Nagaraj et al., 2025) measure influence by correlating per-sample loss trajectories with validation set dynamics, though at the cost of reduced selection quality compared to gradient-based methods. Training-Dynamics-Based Selection methods identify informative examples without an explicit target set. Approaches like TAGCOS (Zhang et al., 2025b) and STAFF (Zhang et al., 2025c) use gradient features from warmup checkpoints or surrogate models. Other efficient, forward-only methods such as S2L use small-model losses to guide selection for larger models (Yang et al., 2024). In contrast, TRIM uses targeted, context-aware token-level selection from a few task samples; outperforming the task-agnostic methods (Table˜1).
Token-Centric Selection. Recognizing the limitations of sample-level metrics, the most recent work has shifted towards token-level signals. This emerging paradigm addresses goals such as improving long-context recall (Chen et al., 2025) or performing general-purpose dataset filtering through complex, multi-stage schemes (Fu et al., 2025). TRIM contributes to this emerging token-centric paradigm but is distinct in its goal and mechanism. Unlike methods for general pruning or long-context analysis, TRIM is specifically built for efficient, targeted selection guided by a small set of samples. It utilizes a novel lightweight saliency heuristic derived from multi-layer attention flow to quantify token importance (Abnar and Zuidema, 2020). By adapting this principle into a single-pass algorithm, TRIM offers a uniquely scalable solution, while outperforming state-of-the-art coreset generation techniques. A more comprehensive review of related works is provided in appendix˜A.
@@ 3 Methodology
We begin with a pretrained language model $\mathbf{M}_{0}$. Our goal is to select a compact coreset $\mathcal{C}$ from a large and diverse instruction-tuning corpus $\mathcal{S}=\{s_{1},\dots,s_{|\mathcal{S}|}\}$, which serves as the candidate pool for fine-tuning $\mathbf{M}_{0}$ on a target domain $\mathcal{T}$. Coreset selection is guided by a small validation set $\mathcal{T}_{\text{val}}$, consisting of as few as 551010 samples from $\mathcal{T}$. Similar to prior works (Xia et al., 2024), we perform a brief warmup phase to adapt $\mathbf{M}_{0}$ to the source distribution and stabilize the attention readouts and token representations used for sample scoring. Specifically, we fine-tune $\mathbf{M}_{0}$ on a small random subset $\mathcal{S}_{\text{warmup}}\subset\mathcal{S}$ (about 5%), which produces the scoring model $\mathbf{M}_{\text{warmup}}$. All subsequent computations, i.e., attention-based saliency, fingerprint construction from $\mathcal{T}_{\text{val}}$, and candidate scoring over $\mathcal{S}$, are performed using only forward passes through $\mathbf{M}_{\text{warmup}}$. TRIM then executes a forward-only, two-stage pipeline guided by $\mathcal{T}_{\text{val}}$. In Stage I (Section˜3.1), it reads attention signals from $\mathbf{M}_{\text{warmup}}$ to construct token-wise, task-defining fingerprints from the examples in $\mathcal{T}_{\text{val}}$. In Stage II (Section˜3.2), each candidate $c\in\mathcal{S}$ is scored by comparing its token representations to the fingerprints, aggregating these comparisons into an example-level score, and selecting the top-ranked examples to form the final coreset $\mathcal{C}$.
@@ 3.1 Stage I: Building Saliency-Weighted Token Fingerprints
Given the target validation set $\mathcal{T}_{\text{val}}$, TRIM constructs a dictionary of token fingerprints. This is achieved through forward passes over $\mathbf{M}_{\text{warmup}}$, where we extract attention signals from the final $L$ layers.
@@
We define a token’s saliency by integrating complementary signals from the row (query) and column (key) dimensions of the attention feature map (Figure˜1a).
Row Saliency. We score a token at position $i$ by the sharpness of its attention distribution. Tokens that focus strongly on a few positions (keys) are considered more salient, while those that spread attention broadly are less informative. We capture this notion of attention sharpness as a proxy for token importance using the entropy of the attention distribution. For layer $l$, head $h$ and sequence length $T$, let $A^{l,h}\in\mathbb{R}^{T\times T}$ denote the attention matrix. We drop $h$ and $l$ notation for simplicity, and compute the entropy for $i^{th}$ row (query) as follows:
@@
Here, $\varepsilon$ is added for numerical stability, and we further normalize this entropy by the number of non-masked keys to obtain the row saliency score:
@@
Column Saliency. The column saliency score is computed by aggregating all attention weights for each column (key) token. This signifies which tokens are attended strongly by others and are hence more crucial for final performance. We quantify this by measuring average attention weights for $j^{th}$ column as follows:
@@
$K^{\text{raw}}_{j}\;=\;\frac{1}{LH}\sum_{l=1}^{L}\sum_{h=1}^{H}k^{l,h}_{j}$ (5)
Min-max scaling is applied to normalize $K^{\text{raw}}_{j}$:
$K_{j}\;=\;\frac{K^{\text{raw}}_{j}-\min K_{j}^{\text{raw}}}{\max K_{j}^{\text{raw}}-\min K_{j}^{\text{raw}}+\varepsilon}\;\in[0,1].$ (6)
Here, $\varepsilon$ ensures numerical stability.
@@
For each sample $v\in\mathcal{T}_{\text{val}}$ in the target validation set, we construct token fingerprints using the aggregated token saliency scores. Fingerprints are defined at the token-class level, where each class corresponds to a unique token in the vocabulary Wu and Papyan (2024). Let $t$ denote a token class, and $h_{v,i}$ be the last layer’s hidden state for token $t$ at position $i$ in a sample $v$. We collect the set of all such occurrences:
@@
The fingerprint for class $t$, denoted $f_{t}$, is then defined as the saliency-weighted average of the normalized hidden states $\hat{h}_{v,i}=\frac{h_{v,i}}{\|h_{v,i}\|_{2}}$:

### v2

@@ Abstract
Instruction tuning is essential for aligning large language models (LLMs) to downstream tasks and commonly relies on large, diverse corpora. However, small, high-quality subsets, known as coresets, can deliver comparable or superior results, though curating them remains challenging. Existing methods often rely on coarse, sample-level signals like gradients, an approach that is computationally expensive and overlooks fine-grained features. To address this, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention), a forward-only, token-centric framework. Instead of using gradients, TRIM operates by matching underlying representational patterns identified via attention-based “fingerprints” from a handful of target samples. Such an approach makes TRIM highly efficient and uniquely sensitive to the structural features that define a task. Coresets selected by our method consistently outperform state-of-the-art baselines by up to 9% on downstream tasks and even surpass the performance of full-data fine-tuning in some settings. By avoiding expensive backward passes, TRIM achieves this at a fraction of the computational cost. These findings establish TRIM as a scalable and efficient alternative for building high-quality instruction-tuning datasets.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have become the de facto standard for a wide range of tasks, but their full potential is unlocked only after a critical, computationally expensive step: instruction tuning, where a pretrained model is fine-tuned to align its outputs with user intent. While large, diverse instruction-tuning corpora have shown to be effective (Ouyang et al., 2022; Chung et al., 2024), a fundamental challenge persists: not all data is created equal. Recent findings have demonstrated that carefully curating a small, high-quality subset, or a “coreset”, can not only match but even surpass the performance of finetuning on the full dataset, all while using a fraction of the compute (Xia et al., 2024; Zhang et al., 2025a).
Current solutions for obtaining coresets typically estimate sample-level importance using influence-style signals based on gradients or Hessians (Xia et al., 2024; Zhang et al., 2025b; San Joaquin et al., 2024). While effective, they are often costly to scale due to per-sample backward passes leading to large compute requirements (Pruthi et al., 2020). Cheaper alternatives based on representation similarity (Hanawa et al., 2021; Zhang et al., 2018) reduce compute by relying on forward-only embeddings, but still assign a single score per sample. It has been shown that sample-level importance estimation leads to two systematic biases: (i) length bias, since aggregating token-level signals (e.g., loss/gradient magnitudes) makes scores depend on sequence length, and gradient norms can be negatively correlated with length in instruction tuning (Xia et al., 2024); and (ii) loss-centric bias, since next-token prediction spreads credit uniformly across tokens even though attention and task-relevant evidence are typically sparse and localized (Zhang et al., 2023; Ge et al., 2023; Lin et al., 2024c; Jiang et al., 2024).
@@
In principle, the most faithful measure of influence would come from token-wise gradients (or Hessians) (Lin et al., 2024a; Chen et al., 2025b). However, this is infeasible at scale since it requires gradient/Hessian information at every token position in a sequence. We therefore aim to develop efficient proxies that preserve token-level resolution while avoiding the prohibitive cost of gradient computation. This leads to a central question: “Given only a few target examples, how can we faithfully and efficiently identify high-impact instruction data by moving from sample-level scoring to token-level signals for task-specific fine-tuning?”
@@
Figure 1: Overview of TRIM. (a) From multi-layer attention, we derive an aggregated token saliency signal that combines row (allocation sharpness) and column (received attention) signals. (b) Using a small target set, we compute one fingerprint per token class by saliency-weighted averaging of last-layer hidden states, capturing task-defining patterns (Section 3.1). (c) For each candidate sample, token states are matched to their class fingerprints (cosine similarity), and token scores are pooled into a single relevance score for ranking (Section 3.2); top-ranked samples form the instruction-tuning coreset.
@@
To address this, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention). Our approach shifts from the traditional sample-level, loss-centric paradigm to a token-level, representation-centric one. The core insight is that a sample’s utility often resides not in the entire sequence but in a few informative tokens (Jiang et al., 2024; Jin et al., 2024); thus, we leverage the model’s attention layers to weight and match these critical representations to a target task.
@@
As shown in Figure 1, TRIM is a two-stage pipeline requiring only forward passes. Stage I constructs lightweight token fingerprints from a small target set, representations that summarize how a given token class appears in the target context. It does so by computing an aggregated token saliency score that combines two complementary signals: row saliency, which measures how sharply a token distributes its own attention, and column saliency, which measures how much attention it receives from other tokens (panel a). For each token class, we then compute a fingerprint as the saliency-weighted mean of the last-layer hidden states of all target-set tokens in that class (panel b), yielding a concise summary of what that class looks like to the model. Stage II uses these fingerprints to score samples in the large candidate corpus. For each sample, we score token-level alignment to the task fingerprint (via cosine similarity; panel c) and pool these scores into a single relevance score. We then select the top-scoring samples to form the final coreset. This design yields three advantages. First, TRIM delivers superior performance and efficiency: with only a handful of samples from the target task, it surpasses state-of-the-art methods (Xia et al., 2024; Zhang et al., 2025b; Yang et al., 2024) while running orders of magnitude faster, owing to its forward-only design that avoids per-sample gradient computation (Sections 4.1 and 5). Second, it offers high structural fidelity by matching core representational patterns (e.g., syntax, mathematical operators), which benefits challenging downstream tasks (Section 4.2). Finally, its token-level scoring inherently mitigates length bias, a common failure mode of sample-level methods (Section 4.4). Beyond targeted selection, we show that TRIM can be adapted to in-domain selection without a separate target set, using only the training pool itself (Section 4.5). Our contributions are:
  • We introduce TRIM, a token-level, representation-based coreset selection method that uses attention-derived token saliency to score and select instruction-tuning examples from only a few target samples.
@@
  • TRIM is a forward-only algorithm that is orders of magnitude faster than gradient-based alternatives, enabling efficient data selection over massive instruction corpora.
@@
  • With only a 5% coreset and a handful of target samples, TRIM outperforms state-of-the-art methods by up to 9% on downstream reasoning tasks, and even exceeds full-data fine-tuning on some tasks.
@@ 2 Related Work
The challenge of curating coresets for instruction tuning has motivated a wide body of work on automated data selection (Cao et al., 2024; Chen et al., 2024a; Zhou et al., 2023; Ding et al., 2023). Existing methods can be grouped into several paradigms. Influence-Based Selection methods estimate data importance by quantifying the causal effect of a candidate example on a target task. As exact influence functions (Koh and Liang, 2017) are infeasible for LLMs, prior work relies on approximations, including Hessian-based variants (Kwon et al., 2024; San Joaquin et al., 2024; Lin et al., 2024b) and first-order alternatives such as checkpointed gradient similarity (e.g., LESS (Xia et al., 2024)). Forward-only proxies such as CLD (Nagaraj et al., 2025), DynUnc (He et al., 2024) correlate per-sample loss trajectories with validation dynamics, trading accuracy for efficiency. Training-dynamics-based methods identify informative examples without an explicit target gradient. Works such as TAGCOS and STAFF use gradient features from warmup checkpoints or surrogate models (Zhang et al., 2025b, c), while S2L transfers small-model loss signals to guide selection for larger models (Yang et al., 2024). In contrast, TRIM uses targeted, context-aware token-level selection from a few task samples. Recognizing the limitations of sample-level, token-centric selection uses token-level signals for pruning, long-context analysis, or filtering (Chen et al., 2025a; Fu et al., 2025). Optimization-based approaches such as QCS (Chen et al., 2025b) cast selection as a bi-level problem coupling sequence selection with token mining, which can be computationally heavy at scale. TRIM contributes to this direction by offering an efficient, reference-guided token-level selector that produces a single relevance score per example for scalable coreset construction. An extended discussion is provided in Appendix A.
@@ 3 Methodology
We begin with a pretrained language model $\mathbf{M}_{0}$. Our goal is to select a compact coreset $\mathcal{C}$ from a large and diverse instruction-tuning corpus $\mathcal{S}=\{s_{1},\dots,s_{|\mathcal{S}|}\}$, which serves as the candidate pool for fine-tuning $\mathbf{M}_{0}$ on a target domain $\mathcal{T}$. Coreset selection is guided by a small validation set $\mathcal{T}_{\text{val}}$ (disjoint from the test set), consisting of as few as 551010 samples from $\mathcal{T}$. Similar to prior works (Xia et al., 2024), we perform a brief warmup phase to adapt $\mathbf{M}_{0}$ to the source distribution. Specifically, we fine-tune $\mathbf{M}_{0}$ on a small random subset $\mathcal{S}_{\text{warmup}}\subset\mathcal{S}$ (about 5%), which produces the scoring model $\mathbf{M}_{\text{warmup}}$. All subsequent computations, i.e., attention-based saliency, fingerprint construction from $\mathcal{T}_{\text{val}}$, and candidate scoring over $\mathcal{S}$, are performed using only forward passes through $\mathbf{M}_{\text{warmup}}$. TRIM then executes a forward-only, two-stage pipeline guided by $\mathcal{T}_{\text{val}}$. In Stage I (Section 3.1), it reads attention signals from $\mathbf{M}_{\text{warmup}}$ to construct token-wise, task-defining fingerprints from the examples in $\mathcal{T}_{\text{val}}$. In Stage II (Section 3.2), each candidate $c\in\mathcal{S}$ is scored by comparing its token representations to the fingerprints, aggregating these comparisons into an example-level score, and selecting the top-ranked examples to form the final coreset $\mathcal{C}$.
@@ 3.1 Stage I: Building Saliency-Weighted Token Fingerprints
Given the target validation set $\mathcal{T}_{\text{val}}$, TRIM builds a compact dictionary of token fingerprints; prototype representations that capture how task-relevant tokens appear in context. This is achieved using only forward passes through $\mathbf{M}_{\text{warmup}}$: we (i) compute a per-token importance weight from attention patterns, (ii) use these weights to aggregate contextual token representations from $\mathcal{T}_{\text{val}}$ into class-level fingerprints, and (iii) down-weight very common tokens (e.g., generic instruction formatting) using a corpus-frequency reweighting.
@@
We define a token’s saliency by integrating complementary signals from the row (query) and column (key) dimensions of the attention feature map (Figure 1a).
Row Saliency. We score a token at position $i$ by the row-wise entropy of its attention distribution. Tokens (query) with low row-wise entropy focus strongly on a few positions (keys) and are considered more salient, while those that spread attention broadly have low entropy and are less informative. For layer $l$, head $h$ and sequence length $T$, let $A^{l,h}\in\mathbb{R}^{T\times T}$ denote the attention matrix. We drop $h$ and $l$ notation for simplicity, and compute the entropy for $i^{th}$ row (query) as:
@@
Here, $\varepsilon$ is added for numerical stability, and we normalize by the number of admissible keys under the attention mask (excluding padding; $j\leq i$ in a causal model) to obtain the row saliency score:
@@
Column Saliency. The column saliency score is computed by aggregating all attention weights for each column (key) token. This signifies which tokens are attended strongly by others and are hence more likely to be informative. We quantify this by measuring average attention weights for $j^{th}$ column as follows:
@@
$\overline{K}_{j}\;=\;\frac{1}{LH}\sum_{l=1}^{L}\sum_{h=1}^{H}k^{l,h}_{j}$ (5)
Since $\overline{K}_{j}$ is unbounded, and its scale can vary across sequences, layers, and heads, we rescale it to $[0,1]$ so it is comparable to $Q_{i}$ before combining the two signals. We therefore apply min–max normalization (with $\varepsilon$ added for numerical stability):
$K_{j}\;=\;\frac{\overline{K}_{j}-\min\overline{K}_{j}}{\max\overline{K}_{j}-\min\overline{K}_{j}+\varepsilon}\;\in[0,1].$ (6)
@@
We now convert the saliency-weighted token signals into fingerprints: for each token class, we build a prototype vector that captures both how the token appears in the target context (via its hidden state) and how important that occurrence is (via its saliency). Fingerprints are defined at the token-class level, where each class corresponds to a unique token in the vocabulary (Wu and Papyan, 2024). Let $t$ denote a token class and let $h_{v,i}$ be the last-layer hidden state for an occurrence of $t$ at position $i$ in a target sample $v\in\mathcal{T}_{\text{val}}$. We collect all such occurrences:
@@
We use last-layer hidden states since they are closest to the model’s prediction head and thus provide a task-relevant token representation while keeping fingerprinting lightweight. The fingerprint for class $t$, denoted $f_{t}$, is then defined as the saliency-weighted average of the normalized hidden states $\hat{h}_{v,i}=\frac{h_{v,i}}{\|h_{v,i}\|_{2}}$:
@@
Frequency reweighting.
Eq. (9) weights token occurrences by attention-derived saliency $\alpha_{v,i}$, but very common tokens (e.g., generic instruction framing) can still dominate the aggregation simply because they appear frequently. To reduce this effect, we apply inverse document frequency (IDF) reweighting by replacing $\alpha_{v,i}$ with $\alpha_{v,i}\cdot\mathrm{IDF}(t)$, where $t=\mathrm{class}(v_{i})$ and $\mathrm{IDF}(t)=\log\!\left(\frac{N+1}{\mathrm{df}(t)+1}\right)$ is computed over the candidate pool $\mathcal{S}$. We analyze the impact of IDF in the ablations (Appendix H).